OpenStudy (anonymous):

if 2, a, b, -54 forms a geometric sequence, find the values of a and b. show work

OpenStudy (anonymous):

a/2 = b/a b/a = -54/b solve for a and b

OpenStudy (anonymous):

so its like -26? @sourwing